Glycerin, C3H8O3, is a nonvolatile nonelectrolyte with a density of 1.26 g/mL at 25oC. Calculate the vapor pressure at 25oC of a solution made by adding 50.0 mL of glycerin to 500.0 mL of water. The vapor pressure of pure water at 25oC is 23.8 torr.
1
Expert's answer
2020-12-14T14:46:01-0500


Partial pressure by solvent (when a non volatile solute is used) = "mole fraction of solvent * vapour pressure of pure solvent"

Therefore, no. of moles of water in 500mL = 500/18

= 27.78

no. of moles of glycerine ="\frac{density * volume}{Molar mass of glycerine}"

=1.26*50/92.1

= 0.684

total moles in solution = 28.464

vapour pressure of solution = "27.78*23.8\/28.64"

= 23.085 Torr